# REST: Diffusion-based Real-time End-to-end Streaming Talking Head Generation via ID-Context Caching and Asynchronous Streaming Distillation

### v1

@@ 3.4 AR-style Streaming Inference
Leveraging the ID-Context Cache and ASD strategies, the streaming model acquires the capability to condition current chunk synthesis on historical generative distributions. Based on this, we formulate an AR-style inference pipeline. The inference pipeline employs a dual-loop inference scheme, where the inner loop performs intra-chunk temporal diffusion and the outer loop executes inter-chunk recursion, progressively propagating historical information and ID anchors through ID-Context Cache. The resulting velocity flow is resolved to the latent space by an FM scheduler and reconstructed to pixel space via the temporal VAE decoder. During inference, we incorporate the Joint CFG (Wang et al., 2025a) mechanism for classifier-free guidance. The complete inference pipeline is detailed in Algorithm 1.
Algorithm 1 Asynchronous Reverse Process
@@
Speech latents $\boldsymbol{E}\in\mathbb{R}^{N\times h_{w}\times d_{a}}$;
@@
Zeros ID-Context Cache $\boldsymbol{C}=\{\boldsymbol{C}_{1},\cdots,\boldsymbol{C}_{n}\}$.
@@
Generated latents $\boldsymbol{Z}(0)\in\mathbb{R}^{h\times w\times N\times d_{v}}$.
for $i\leftarrow 1$ to $k$ do
@@
[…]
$\boldsymbol{Z}_{j}(T_{i+1}),\boldsymbol{C}_{j+1}\xleftarrow[\text{FM}]{\text{CFG}}\mathcal{S}_{\theta}\big(\boldsymbol{Z}_{j}(T_{i}),\boldsymbol{E}_{j},\boldsymbol{t}_{j},\boldsymbol{C}_{j}\big)$
@@ 4.1 Experimental Setup
Implementation Details. Experiments encompassing both training and inference are conducted on HDTF (Zhang et al., 2021) and MEAD (Wang et al., 2020) datasets. 95% data of both datasets is randomly allocated for training and the remaining 5% for testing, ensuring no overlap between the partitions. We employ a two-stage training strategy. In the first stage, the non-streaming teacher model is pre-trained at a resolution of $512\times 512$ pixels with a learning rate of $1\times 10^{-5}$. In the second stage, the streaming student model with ID-Context Cache is trained at a resolution of $512\times 512$ pixels and 121 frames, with a learning rate of $1\times 10^{-5}$ and a batch size of 1. All reported results use 8-step sampling and Joint-CFG with $\alpha=6.0$ unless specified. Both training and evaluation are performed on NVIDIA A100 GPUs.
Evaluation Metrics. Generation performance is assessed using several metrics. For visual quality, we employ the Fréchet Inception Distance (FID) (Seitzer, 2020) for image-level realism between synthesized videos and reference images and the Fréchet Video Distance (FVD) (Unterthiner et al., 2019) for frame-level realism between synthesized and ground-truth videos; lower values indicate better performance for both metrics. Lip synchronization is measured with SyncNet (Chung & Zisserman, 2017), where a higher Synchronization Confidence (Sync-C) and a lower Synchronization Distance (Sync-D) indicate superior alignment with speech input. We further use the Expression-FID (E-FID) metric from EMO (Tian et al., 2024) to measure the expression divergence between synthesized and ground-truth videos, with lower values indicating more faithful reproduction of expressions. Finally, we evaluate the efficiency of the diffusion backbone of each model by measuring the average runtime of the backbone per video (Runtime).
@@ 4.3 Ablation Studies
Ablation Studies on ID-Context Cache. To investigate the specific contributions of the ID-Sink and Context-Cache principles within our proposed ID-Context Cache methodology, we conduct an ablation study under different experimental configurations, as follows:
@@
  • w/o ID-Sink: Configuration exclude the ID‑Sink principle, retaining only the Context‑Cache principle.
  • w/o Context-Cache: Configuration exclude the ID-Sink principle, retaining only the ID-Sink principle.
The inference segment length is set to five times the training chunk length (120 frames). The performance of each configuration is evaluated on the HDTF dataset across FID, FVD, and SyncNet metrics. As detailed in Tab. 2, ablating the ID-Sink mechanism has a minimal impact on Sync-C but significantly worsens FID and FVD, indicating that suboptimal ID consistency leads to semantic distortion in generated videos. In contrast, ablating the Context-Cache mechanism results in a clear deterioration in Sync-C, accompanied by declines in FID and FVD, which can be attributed to weakened temporal consistency. To further investigate these effects, we conducted a visual analysis by uniformly sampling frames at aligned timestamps and computing difference heatmaps between adjacent frames. The results are shown in Fig. 3.
As observed, the generation results of w/o ID‑Sink configuration exhibit progressively deteriorating ID consistency over time, accompanied by a degradation in fine-grained details (e.g., the eyes in the zoomed-in views). Moreover, the overall color tone gradually degrades as inference proceeds, and the error heatmap shows high‑error regions not only on the subject but also across the background, indicating suboptimal ID consistency. Conversely, the w/o Context-Cache configuration suffers from compromised motion smoothness, characterized by abrupt motion discontinuities at chunk boundaries (see zoomed-in views). In addition, the difference heatmap displays overall higher error values distributed across the entire figure, confirming inferior temporal consistency and motion smoothness. In contrast, the Full ID-Context Cache, which integrates both principles, achieves superior performance in both identity preservation and motion smoothness. The zoom‑in details are clearly superior to those of the ablated variants. Notably, in the difference heatmaps of the results of Full ID‑Context Cache, high values are concentrated solely on dynamic facial regions (i.e., eyes and mouth), while the background and overall head structure exhibit minimal error. These visualizations confirm the crucial roles of ID-Sink and Context-Cache in enhancing identity consistency and temporal smoothness, demonstrating the effectiveness of our proposed method.
Ablation Studies on ASD. To validate the effectiveness of our proposed ASD strategy and the contribution of each loss terms, we conduct a quantitative ablation study using the following configurations:
@@
We report the FID, FVD, and SyncNet scores of each configuration on the HDTF dataset, as shown in Tab 3. As indicated by the results, removing the Smooth loss while retaining the Contrastive loss for teacher-student distillation leads to a deterioration in the FVD metric, whereas FID and Sync-C scores show only marginal degradation. This indicates that the Smooth loss plays a crucial role in enhancing temporal coherence and preserving the overall semantic consistency of the generated videos. Conversely, removing the Contrastive loss and retaining only the Smooth loss results in noticeable deterioration in both Sync-C and FID metrics, with minimal impact on FVD. This demonstrates that the Contrastive loss primarily contributes to improving the overall visual quality and audio-visual alignment of the student model during streaming generation. Furthermore, the complete removal of the ASD strategy shows a substantial performance drop across all three metrics compared to the Full ASD configuration. These results validate the significant efficacy of our non-streaming to streaming distillation strategy in improving both generation quality and temporal consistency.
@@ 4.4 Case Study
To provide a more intuitive comparison of the generation quality among SOTA THG methods, we present a visual analysis of a representative test case. Conditioned on the identical reference image and driving audio, we utilize each model to generate talking head videos of equal length of 120 frames. Key frames at identical timestamps are sampled for visual comparison, as illustrated in Fig. 4, with the inference runtime of backbone network for each model provided.

### v2

@@ 3.4 AR-style Streaming Inference
Leveraging the ID-Context Cache and ASD strategies, we formulate an AR-style inference pipeline. The pipeline employs a dual-loop inference scheme, where the inner loop performs intra-chunk diffusion and the outer loop executes inter-chunk recursion, progressively propagating historical information and ID anchors through ID-Context Cache. The resulting velocity flow is resolved to the latent space by an FM scheduler and reconstructed to pixel space via the temporal VAE decoder. During inference, we incorporate the Joint CFG (Wang et al., 2025a) mechanism for classifier-free guidance. The complete pipeline is detailed in Algorithm 1.
Algorithm 1 AR-style Streaming Inference
@@
Speech latents $\boldsymbol{E}\in\mathbb{R}^{L\times h_{w}\times d_{a}}$, chunk length $l$;
@@
Zeros ID-Context $\boldsymbol{C}=\{\boldsymbol{C}_{1},\cdots,\boldsymbol{C}_{k}\}$ ($\boldsymbol{C}_{0}=\emptyset$).
@@
Generated latents $\boldsymbol{Z}(0)\!\in\!\mathbb{R}^{h\times w\times L\times d_{\text{v}}}$ ($L$ frames).
$\boldsymbol{Z}(T_{1})\leftarrow\{\boldsymbol{Z}_{1}(T_{1}),\cdots,\boldsymbol{Z}_{k}(T_{1})\}$
@@
$\boldsymbol{Z}_{i}(T_{j+1}),\boldsymbol{C}_{i}\xleftarrow[\text{FM}]{\text{CFG}}\mathcal{S}_{\theta}\big(\boldsymbol{Z}_{i}(T_{j}),\boldsymbol{E}_{i},\boldsymbol{C}_{i-1},\boldsymbol{t}_{j}\big)$
@@ 4.1 Experimental Setup
Implementation Details. Experiments encompassing both training and inference are conducted on HDTF (Zhang et al., 2021) and MEAD (Wang et al., 2020) datasets. 95% data of both datasets is randomly allocated for training and the remaining 5% for testing, ensuring identity-disjoint of dataset splitting. We employ a two-stage training strategy. In the first stage, the non-streaming teacher model is pre-trained at a resolution of $512\times 512$ pixels with a learning rate of $1\times 10^{-5}$. In the second stage, the streaming student model with ID-Context Cache is trained at a resolution of $512\times 512$ pixels and 97 frames, with a learning rate of $1\times 10^{-5}$ and a batch size of 1. All reported results use 8-step sampling and Joint-CFG with $\alpha=6.0$ unless specified. Both training and evaluation are performed on NVIDIA A100 GPUs.
Evaluation Metrics. Generation performance is assessed using several metrics. For visual quality, we employ the Fréchet Inception Distance (FID) (Seitzer, 2020) for image-level realism between synthesized videos and reference images and the Fréchet Video Distance (FVD) (Unterthiner et al., 2019) for frame-level realism between synthesized and ground-truth videos; lower values indicate better performance for both metrics. Lip synchronization is measured with SyncNet (Chung and Zisserman, 2017), where a higher Synchronization Confidence (Sync-C) and a lower Synchronization Distance (Sync-D) indicate superior alignment with speech input. We further use the Expression-FID (E-FID) metric from EMO (Tian et al., 2024) to measure the expression divergence between synthesized and ground-truth videos, with lower values indicating more faithful reproduction of expressions. Finally, we evaluate the efficiency of the diffusion backbone of each model by measuring the average runtime of the backbone per video (Runtime).
@@ 4.3 Ablation Study
Ablation Study on ID-Context Cache. To investigate the specific contributions of the ID-Sink and Context-Cache principles within our proposed ID-Context Cache methodology, we conduct an ablation study under different experimental configurations, as follows:
@@
  • w/o ID-Sink: Configuration excludes the ID‑Sink principle, retaining only the Context‑Cache principle.
  • w/o Context-Cache: Configuration excludes the Context‑Cache, retaining only the ID-Sink principle.
The inference segment length is set to five times the training chunk length (121 frames). The performance of each configuration is evaluated on the HDTF dataset across FID, FVD, and SyncNet metrics. As detailed in Tab. 2, ablating the ID-Sink mechanism has a minimal impact on Sync-C but significantly worsens FID and FVD, indicating that suboptimal ID consistency leads to semantic distortion in generated videos. In contrast, ablating the Context-Cache mechanism results in a clear deterioration in Sync-C, accompanied by declines in FID and FVD, which can be attributed to weakened temporal consistency. To further investigate these effects, we conducted a visual analysis by uniformly sampling frames at aligned timestamps and computing difference heatmaps between adjacent frames. The results are shown in Fig. 3, where the inference length is 121 frames.
As observed, the generation results of w/o ID‑Sink configuration exhibit progressively deteriorating ID consistency over time, accompanied by a degradation in fine-grained details (e.g., the eyes in the zoomed-in views). Moreover, the overall color tone gradually degrades as inference proceeds, and the error heatmap shows high‑error regions not only on the subject but also across the background, indicating suboptimal ID consistency. Conversely, the w/o Context-Cache configuration suffers from compromised motion smoothness, characterized by abrupt motion discontinuities at chunk boundaries (see zoomed-in views). In addition, the difference heatmap displays overall higher error values distributed across the entire figure, confirming inferior temporal consistency and motion smoothness. In contrast, the Full ID-Context Cache, which integrates both principles, achieves superior performance in both identity preservation and motion smoothness. The zoom‑in details are clearly superior to those of the ablated variants. Notably, in the difference heatmaps of the results of Full ID‑Context Cache, high values are concentrated solely on dynamic facial regions (e.g., eyes and mouth), while the background and overall head structure exhibit minimal error. These visualizations confirm the crucial roles of ID-Sink and Context-Cache in enhancing identity consistency and temporal smoothness, demonstrating the effectiveness of our proposed method.
Ablation Study on ASD. To validate the effectiveness of our proposed ASD strategy and the contribution of each loss terms, we conduct a quantitative ablation study using the following configurations:
@@
We report the FID, FVD, and SyncNet scores of each configuration on the HDTF dataset, as shown in Tab 3. As indicated by the results, removing the Smooth loss while retaining the Contrastive loss for teacher-student distillation leads to a deterioration in the FVD metric, whereas FID and Sync-C scores show only marginal degradation. This indicates that the Smooth loss plays a crucial role in enhancing temporal coherence and preserving the overall semantic consistency of the generated videos. Conversely, removing the Contrastive loss and retaining only the Smooth loss results in noticeable deterioration in both Sync-C and FID metrics, with minimal impact on FVD. This demonstrates that the Contrastive loss primarily contributes to improving the overall visual quality and audio-visual alignment of the student model during streaming generation. Furthermore, the complete removal of the ASD strategy shows a substantial performance drop across all three metrics compared to the Full ASD configuration. These results validate the significant efficacy of our non-streaming to streaming distillation strategy in improving both video quality and temporal consistency of semi-autoregressive streaming generation.
@@ 4.4 Case Study
To provide a more intuitive comparison of the generation quality among SOTA THG methods, we present a visual analysis of a representative test case. Conditioned on the identical reference image and driving audio, we utilize each model to generate talking head videos of equal length of 121 frames. Key frames at identical timestamps are sampled for visual comparison, as illustrated in Fig. 4, with the inference runtime of backbone network for each model provided.